OpenStudy (anonymous):

super easy but i'm stuck with it ! solve x^3-x^2+2=0 ...

Directrix (directrix):

x^3-x^2+2=0 (x + 1) (x^2 -2x +2) = 0 Using the Zero Product Property x + 1 = 0 or x^2 - 2x + 2 = 0 and ..

OpenStudy (anonymous):

how did you factor the first step ? can you explain ? =]

Directrix (directrix):

x = -1 or x^2 - 2x + 2 = 0 To get the roots of the quadratic, use the Quadratic Formula or Technology. Either way, you will get those two roots to be the cojugate pair (1 + i) and (1 - i). The real root is -1.

Directrix (directrix):

I used the Rational Root Theorem to get candidates for rational roots. I tried -1 first. Then, the cubic was reduced to the quadratic with the complex roots.

OpenStudy (anonymous):

ok thank you very much !

OpenStudy (anonymous):

thanks again ! =]

Directrix (directrix):

Read Descartes' Rule of Signs to use with the Rational Root Theorem: http://www.purplemath.com/modules/drofsign.htm